OpenStudy (anonymous):

Simplify: 3(n+1)! / 5n!

OpenStudy (anonymous):

= 3(n+1)n!/5n! = 3(n+1)/5

OpenStudy (anonymous):

How did you get taht answer?

OpenStudy (anonymous):

do the n! just cancel out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because (n+1)!= (n+1)n! , and n!= n(n-1)(n-2)! , ... etc.